# Mario Shannon

### v1

@@ 2.1. Anosov flows on sol-manifolds
There exists a homeomorphism $\tilde{M}\to\mathbb{R}^{3}$ taking $\tilde{\mathcal{O}}$ to the foliation by vertical lines $\mathcal{V}=\{z\times\mathbb{R}:z\in\mathbb{R}^{2}\}$ of $\ \mathbb{R}^{3}=\mathbb{R}^{2}\times\mathbb{R}$.
As a consequence, the quotient $\Omega\coloneqq\tilde{M}/\tilde{\mathcal{O}}$ is homeomorphic to $\mathbb{R}^{2}$. The center-stable and center-unstable foliations of $M$ lift to a pair of codimension $=1$ transverse foliations of $\tilde{M}$ that intersect along $\tilde{\mathcal{O}}$-leaves, and hence induce a pair of 1-dimensional transverse foliations of $\Omega$. We denote them by $\mathcal{F}^{s}_{\Omega}$ and $\mathcal{F}^{u}_{\Omega}$.
The triple $\left(\Omega;\mathcal{F}^{s}_{\Omega},\mathcal{F}^{u}_{\Omega}\right)$ is called the bi-foliated plane associated to $(\phi_{t},M)$. A general bi-foliated plane may have several different natures. A case of particular interest is given by the following result:
@@
Consider the quotient spaces $\Lambda^{s}\coloneqq\Omega/\mathcal{F}^{s}_{\Omega}$ and $\Lambda^{u}\coloneqq\Omega/\mathcal{F}^{u}_{\Omega}$. Then $\Lambda^{s}$ is homeomorphic to the real line iff $\Lambda^{u}$ is homeomorphic to the real line, and in this case there are two possibilities, represented in figure 2:
@@
    Either $(\Omega;\mathcal{F}^{s}_{\Omega},\mathcal{F}^{u}_{\Omega})$ is equivalent to $\mathbb{R}^{2}$ endowed with the foliations by horizontal and vertical lines. In this case $(\phi_{t},M)$ is orbitally equivalent to the suspension flow generated by a hyperbolic matrix $A\in\mathrm{SL}(2,\mathbb{Z})$.
@@
    Or $(\Omega;\mathcal{F}^{s}_{\Omega},\mathcal{F}^{u}_{\Omega})$ is equivalent to the space obtained by restricting the horizontal and vertical foliations of $\mathbb{R}^{2}$ onto an open band delimited by two parallel lines, none of angle $\theta=0$ or $\theta=\pi/2$.
When $\Lambda^{s}\simeq\Lambda^{u}\simeq\mathbb{R}$ the flow $(\phi_{t},M)$ is said to be $\mathbb{R}$-covered. In case (i) the bi-foliated plane is said to have global product structure, while in case (ii) is said to be skewed $\mathbb{R}$-covered, and it can be positively or negatively twisted according to the sign of $-\pi/2<\theta<\pi/2$. Case (i) is rigid, in the sense that there is only one family of Anosov flows admitting this bi-foliated plane structure, namely, the suspension Anosov flows. Case (ii) is more flexible, there are several families of Anosov flows having this kind of bi-foliated plane. For instance, geodesic flows on hyperbolic surfaces and contact Anosov flow are in this category (see e.g. [5]). We will use the following result:
Proposition 16 (See [11]. See also [3]).
@@
Theorem 17 (Barthelme-Fenley, [1]).
If $\psi_{t}:N\to N$ is a $\mathbb{R}$-covered Anosov flow on a hyperbolic 3-manifold, then for every periodic orbit $\delta$ there exists an infinite set $\mathcal{FH}(\delta)$ of periodic orbits which are equivalent to $\delta$ as knots. That is, for every $\beta\in\mathcal{FH}(\gamma)$ there exists a 1-parameter family of homeomorphisms $H_{t}:M\to M$ such that $H_{0}=\textrm{id}_{M}$ and $H_{1}(\beta)=\delta$ preserving orientation of these curves.
@@ 2.4. Birkhoff sections
Theorem A can be rephrased in terms of Birkhoff sections. Given an Anosov flow $\phi_{t}:M\to M$, a Birkhoff section is an immersion of a compact surface $\Sigma$ inside $M$, denoted $\iota:(\Sigma,\partial\Sigma)\to(M,\Gamma)$, satisfying that:
@@
    On the boundary $\partial\Sigma$ the map $\iota$ is a covering onto a finite union of $\phi_{t}$-periodic orbits $\Gamma=\{\gamma_{1},\dots,\gamma_{k}\}$;
@@
    On the interior $\mathring{\Sigma}\coloneqq\Sigma\setminus\partial\Sigma$ the map $\iota$ is an embedding and $\iota(\mathring{\Sigma})$ is transverse to the $\phi_{t}$-orbits;
@@
For each $\gamma_{i}\in\Gamma$ we will set:
@@
  • $m_{i}=m(\gamma_{i},\Sigma)\in\mathbb{Z}$ to be the degree of the covering $\iota:C_{ij}\to\gamma_{i}$, where $C_{i1},\dots,C_{ip_{i}}$ are the components of $\iota^{-1}(\gamma_{i})$. This integer is called the multiplicity of the section $\Sigma$ at $\gamma_{i}$ and is the same for every component (see e.g. [16]).
Properties (1), (2), (3) imply that there exists a first return map $P_{\Sigma}:\mathring{\Sigma}\to\mathring{\Sigma}$ on the interior of the section, and hence the restriction of $\phi_{t}$ to $M\setminus\Gamma$ is a suspension flow (in an open manifold). There is a construction called blow-down which consists collapsing each component of $\partial\Sigma$ into a point. This produces a closed surface $\hat{\Sigma}$ together with a homeomorphism $\hat{P}_{\Sigma}:\hat{\Sigma}\to\hat{\Sigma}$ induced from $P_{\Sigma}$. For each $\gamma_{i}\in\Gamma$, the set of components $\iota^{-1}(\gamma_{i})$ corresponds to a periodic orbit of $\hat{P}_{\Sigma}$ with period $=p_{i}$. The feature of this construction is that $(\phi_{t},M)$ can be obtained from the suspension flow $(\hat{\phi}_{t},\hat{M})$ generated by the blow-down homeomorphism, by making Fried surgeries on the curves $\gamma_{1},\dots,\gamma_{k}$. See [12] for more details.
If $\Sigma$ has genus $=1$, since $\phi_{t}:M\to M$ is Anosov and $\mathring{\Sigma}$ is transverse to the $\phi_{t}$-orbits, then the $cs/cu$-foliations print on $\mathring{\Sigma}$ a pair of transverse foliations $\mathcal{F}^{s}_{\Sigma}$, $\mathcal{F}^{u}_{\Sigma}$, that extend onto foliations on $\hat{\Sigma}$ (without singularities, since $\chi(\hat{\Sigma})=0$), that are respectively contracted and expanded by the action of $\hat{P}_{\Sigma}$. Thus, the homeomorphism $\hat{P}:\hat{\Sigma}\to\hat{\Sigma}$ is conjugated to a linear Anosov diffeomorphism on the torus given by a matrix $A\in\mathrm{SL}(2,\mathbb{Z})$. The Birkhoff section actually exhibits a path of Fried surgeries connecting $(\phi_{t},M)$ with the suspension flow $(\phi_{t}^{A},M_{A})$:
$(\phi_{t},M)\xrightarrow[]{(\gamma_{1},-m_{1}),\dots,(\gamma_{k},-m_{k})}(\phi_{t}^{A},M_{A}).$
@@
Given a periodic orbit $\alpha$ of the suspension flow $(\phi_{t}^{A},M_{A})$ define $\textsl{per}_{\mathbb{Z}}(\alpha)$ to equal the pairing $[\alpha]\cdot[\mathbb{T}^{2}]$ of the homology class $[\alpha]\in H_{1}(M_{A})$ with a generator $[\mathbb{T}^{2}]$ of $H^{1}(M_{A};\mathbb{Z})$.
@@
Let $A\in\mathrm{SL}(2,\mathbb{Z})$ be a matrix with $\textsc{tr}(A)\geq 3$ satisfying that $A$ is conjugated to $A^{-1}$ in $\mathrm{GL}(2,\mathbb{Z})$. Then, given any periodic orbit $\gamma$ of the suspension flow $(\phi_{t}^{A},M_{A})$ there exists $m_{0}=m_{0}(\gamma)\geq 1$ such that, for every integer $m$ with $|m|\geq m_{0}$ there are $\infty$-many Birkhoff sections $\iota:(\Sigma,\partial\Sigma)\to(M_{A},\Gamma)$ verifying that:
@@
      $p(\gamma,\Sigma)=\textsl{per}_{\mathbb{Z}}(\alpha)$, $p(\alpha,\Sigma)=\textsl{per}_{\mathbb{Z}}(\gamma)$,
@@ 2.5. Construction of the paths
Let’s prove Theorem A-A’ and Proposition B. For this, fix a matrix $A\in\mathrm{SL}(2,\mathbb{Z})$ with $\textsc{tr}(A)\geq 3$, choose some periodic orbit $\gamma$ of the suspension flow $\phi_{t}^{A}:M_{A}\to M_{A}$, let $m>0$ be an integer and define $(\psi_{t},N)$ to be the flow obtained from the suspension $(\phi_{t}^{A},M_{A})$ by doing a Fried surgery along $\gamma$ with slope $=m$. We denote by $\mathcal{P}:\textsl{Per}(\phi_{t}^{A},M_{A})\to\textsl{Per}(\psi_{t},N)$ the natural bijection between periodic orbits induced by this surgery, and denote $\delta=\mathcal{P}(\gamma)$.
Figure 3. Surgery $(\phi_{t},M)\xrightarrow[]{(\gamma,m),(\alpha,-m)}(\phi_{t}^{B},M_{A})$, $B\simeq A$ or $B\simeq A^{-1}$.
@@
  • By proposition 16, for every $m>0$ then $(\psi_{t},N)$ is a skewed $\mathbb{R}$-covered Anosov flow, positively twisted;
@@
Setting $m\geq m_{0}$ we get a skewed $\mathbb{R}$-covered Anosov flow $\psi_{t}:N\to N$ on a hyperbolic 3-manifold. By theorem 17, there exists an infinite set $\mathcal{FH}(\delta)$ of $\psi_{t}$-periodic orbits, each of them orientation preserving isotopic to $\delta$.
Consider one orbit $\beta\in\mathcal{FH}(\delta)$ different from $\delta$. Since $\delta$ and $\beta$ are equivalent as oriented knots, in particular there exists a homeomorphism on $N$ taking $\delta\mapsto\beta$ and preserving their orientations, so the surgery on $(\psi_{t},N)$ with parameters $(\beta,-m)$ will produce an Anosov flow on a manifold homeomorphic to $M_{A}$. Recall that $M_{A}$ is a sol-manifold, therefore from the classification stated in theorem 9 and proposition 11 this flow must be orbitally equivalent to the suspension of either $A$ or $A^{-1}$. That is:
$(\phi_{t}^{A},M_{A})\xmapsto{(\gamma,m)}(\psi_{t},N)\xmapsto{(\beta,-m)}(\phi_{t}^{B},M_{A}),\ \text{where}\ B\simeq A\ \text{or}\ B\simeq A^{-1}.$
If we assume in addition that $A$ is conjugated to $A^{-1}$ in $GL(2,\mathbb{Z})$ then the surgery described above is a closed loop in $\mathbf{G}_{\circ}$ based at $(\phi_{t}^{A},M_{A})$. Define
  • $\alpha$ to be the preimage of $\beta$ under the surgery $(\phi_{t}^{A},M_{A})\xmapsto{(\gamma,m)}(\psi_{t},N)$,
  • $\eta$ to be the image of $\delta$ under the surgery $(\psi_{t},N)\xmapsto{(\beta,-m)}(\phi_{t}^{A},M_{A})$.
Let $\gamma^{\prime}$ be the image of $\beta$ under the surgery $(\psi_{t},N)\xmapsto{(\beta,-m)}(\phi_{t}^{A},M_{A})$. Along the construction there is a sequence of homeomorphisms $M_{A}\setminus\gamma\to N\setminus\delta\to N\setminus\beta\to M_{A}\setminus\gamma^{\prime}$ so by proposition 12 we have that, up to a self-orbital equivalence of ($\phi_{t}^{A},M_{A})$, the orbit $\gamma^{\prime}$ is equal to either $\gamma$. We have codified paths:
(1) $(\phi_{t}^{A},M_{A})\xmapsto{(\gamma,m),(\eta,-m)}(\phi_{t}^{A},M_{A})\ \text{taking}\ \mathcal{P}:\eta\mapsto\gamma\ \text{and}\ \mathcal{P}:\gamma\mapsto\alpha$
(2) $(\phi_{t}^{A},M_{A})\xmapsto{(\gamma,m),(\alpha,-m)}(\phi_{t}^{A},M_{A})\ \text{taking}\ \mathcal{P}:\alpha\mapsto\gamma\ \text{and}\ \mathcal{P}:\gamma\mapsto\eta.$
The global transverse section of $\phi_{t}^{A}:M_{A}\to M_{A}$ is transformed by the surgery (1) into the desired Birkhoff section $\iota:(\Sigma,\partial\Sigma)\to(M_{A},\Gamma)$ for $(\phi_{t}^{A},M_{A})$, where $\Gamma=\{\gamma,\alpha\}$. In fact:
@@
  • It is direct that $\Sigma$ has $p(\alpha,\Sigma)=\textsl{per}_{\mathbb{Z}}(\gamma)$ boundary components along $\alpha$;
  • To determine $p(\gamma,\Sigma)$, since $[\iota(\partial\Sigma)]=0\in H_{1}(M_{A})\simeq\mathbb{Z}$ we have the relation
    $0=p(\gamma,\Sigma)\cdot\textsl{per}_{\mathbb{Z}}(\gamma)\cdot m(\gamma,\Sigma)+p(\alpha,\Sigma)\cdot\textsl{per}_{\mathbb{Z}}(\alpha)\cdot m(\alpha,\Sigma),$
    from where we conclude that $p(\gamma,\Sigma)=\textsl{per}_{\mathbb{Z}}(\alpha)$.
Finally, the set $\mathcal{Q}(m,\gamma)$ consists in all periodic orbits $\alpha$ of $(\phi_{t}^{A},M_{A})$ such that $\beta=\mathcal{P}(\alpha)$ belongs to $\mathcal{FH}(\mathcal{P}(\gamma))$.
@@
Proposition B follows from an estimation on the number of periodic orbits $\beta$ of length $\leq t$ in $\mathcal{FH}(\delta)$ given in [2].
@@
If $\psi_{t}:N\to N$ is a $\mathbb{R}$-covered Anosov flow in a hyperbolic manifold then there exist constants $C_{0},t_{0}>0$ such that, for every $\psi_{t}$-periodic orbit $\delta$ it is verified that
(3) |$|\{\beta\in\mathcal{FH}(\delta):\textsl{per}(\beta)\leq t\}|\leq C_{0}\sqrt{t}\log(t),\ \forall\ t>t_{0}.$
@@
To prove proposition B it is enough to show that there is $C_{1},t_{1}>0$ such that
(4) |$|\{\alpha\in\mathcal{Q}(\gamma,m):\textsl{per}(\alpha)\leq t\}|\leq C_{1}\sqrt{t}\log(t),\ \forall\ t>t_{1}.$
Because, in this case, it follows that for every $t>t_{1}$ then
$|\mathcal{Q}^{2}_{t}(m)|=|\{(\gamma_{1},\gamma_{2})\in\mathbb{P}_{t}^{2}:\gamma_{2}\in\mathcal{Q}(\gamma_{1},m)\}|\leq C_{1}\sqrt{t}\log(t)|\mathbb{P}_{t}|,$
from where we conclude that
$\lim_{t\to\infty}\frac{|\mathcal{Q}^{2}_{t}(m)|}{|\mathbb{P}^{2}_{t}|}\leq\lim_{t\to\infty}\frac{C_{1}\sqrt{t}\log(t)|\mathbb{P}_{t}|}{|\mathbb{P}_{t}|^{2}}=\lim_{t\to\infty}\frac{C_{1}\sqrt{t}\log(t)}{|\mathbb{P}_{t}|}=0.$
Here we use the fact that $|\mathbb{P}_{t}|\to+\infty$ with exponential order (Bowen-Margulis, see e.g. [13]).

### v2

@@ 2.3. $\mathbb{R}$-covered Anosov flows on hyperbolic 3-manifolds
[…]
As a consequence, the quotient $\Omega\coloneqq\tilde{M}/\tilde{\mathcal{O}}$ is homeomorphic to $\mathbb{R}^{2}$. The center-stable and center-unstable foliations of $M$ lift to a pair of codimension 1 transverse foliations of $\tilde{M}$ that intersect along $\tilde{\mathcal{O}}-$leaves, and hence induce a pair of 1-dimensional transverse foliations of $\Omega$. We denote them by $\mathcal{F}^{\mathrm{s}}_{\Omega}$ and $\mathcal{F}^{\mathrm{u}}_{\Omega}$.
The triple $\left(\Omega;\mathcal{F}^{\mathrm{s}}_{\Omega},\mathcal{F}^{\mathrm{u}}_{\Omega}\right)$ is called the bi-foliated plane associated to $(\phi,M)$. A general bi-foliated plane may have several different natures. A case of particular interest is given by the following result:
@@
Consider the quotient spaces $\Lambda^{s}\coloneqq\Omega/\mathcal{F}^{\mathrm{s}}_{\Omega}$ and $\Lambda^{u}\coloneqq\Omega/\mathcal{F}^{\mathrm{u}}_{\Omega}$. Then $\Lambda^{s}$ is homeomorphic to the real line if and only if $\Lambda^{u}$ is homeomorphic to the real line, and in this case there are two possibilities, represented in Figure (2):
@@
    Either $(\Omega;\mathcal{F}^{\mathrm{s}}_{\Omega},\mathcal{F}^{\mathrm{u}}_{\Omega})$ is equivalent to $\mathbb{R}^{2}$ endowed with the foliations by horizontal and vertical lines. In this case $(\phi,M)$ is orbitally equivalent to the suspension flow generated by a hyperbolic matrix $A\in\mathrm{SL}(2,\mathbb{Z})$.
@@
    Or $(\Omega;\mathcal{F}^{\mathrm{s}}_{\Omega},\mathcal{F}^{\mathrm{u}}_{\Omega})$ is equivalent to the space obtained by restricting the horizontal and vertical foliations of $\mathbb{R}^{2}$ onto an open band delimited by two parallel lines, none of angle $\theta=0$ or $\theta=\pi/2$.
When $\Lambda^{s}\simeq\Lambda^{u}\simeq\mathbb{R}$ the flow $(\phi,M)$ is said to be $\mathbb{R}$-covered. In case (i) the bi-foliated plane is said to have global product structure, while in case (ii) is said to be skewed $\mathbb{R}$-covered, and it can be positively or negatively twisted according to the sign of $\pi/2<\theta<\pi/2$. Case (i) is rigid, in the sense that there is only one family of Anosov flows admitting this bi-foliated plane structure, namely, the suspension Anosov flows. Case (ii) is more flexible, there are several families of Anosov flows having this kind of bi-foliated plane. For instance, geodesic flows on hyperbolic surfaces and contact Anosov flow are in this category (see e.g. [5]). We will use the following result:
Proposition 18 (See [11] and [3]).
@@
This results follows directly from Theorem A at [3], since a Fried surgery supported in one orbit of a suspension Anosov flow will produce an Anosov flow endowed with a positive Birkhoff section, as defined therein.
@@
Theorem 19 (Barthelmé-Fenley, [1]).
If $\{\psi_{t}:N\to N\}_{t\in\mathbb{R}}$ is a $\mathbb{R}$-covered Anosov flow on a hyperbolic 3-manifold, then for every periodic orbit $\delta$ there exists an infinite set $\mathcal{FH}(\delta)$ of periodic orbits which are equivalent to $\delta$ as knots. That is, for every $\beta\in\mathcal{FH}(\gamma)$ there exists a 1-parameter family of homeomorphisms $H_{t}:M\to M$ such that $H_{0}=\textrm{id}_{M}$ and $H_{1}(\beta)=\delta$ preserving orientation of these curves.
@@ 2.4. Birkhoff sections
Theorem A can be rephrased in terms of Birkhoff sections. Given an Anosov flow $\{\phi_{t}:M\to M\}_{t\in\mathbb{R}}$ a Birkhoff section is an immersion of a compact surface $\Sigma$ inside $M$, denoted $\iota:(\Sigma,\partial\Sigma)\to(M,\Gamma)$, satisfying that:
@@
    On the boundary $\partial\Sigma$ the map $\iota$ is a covering onto a finite union of $\phi$-periodic orbits $\Gamma=\{\gamma_{1},\dots,\gamma_{k}\}$;
@@
    On the interior $\mathring{\Sigma}\coloneqq\Sigma\setminus\partial\Sigma$ the map $\iota$ is an embedding and $\iota(\mathring{\Sigma})$ is transverse to the $\phi$-orbits;
@@
Observe that, since $\mathring{\Sigma}$ is transverse to the $\phi$-orbits, there is always a canonical choice of orientation on $\Sigma$, defined by the convention that the orientation of $\Sigma$ paired with the coorientation determined by the flow is equivalent to the opposite orientation of $M$. For each $\gamma_{i}\in\Gamma$ we will set:
@@
  • $m_{i}=m(\gamma_{i},\Sigma)\in\mathbb{Z}$ to be the degree of the oriented covering $\iota:C_{ij}\to\gamma_{i}$, where $C_{i1},\dots,C_{ip_{i}}$ are the components of $\iota^{-1}(\gamma_{i})$ oriented as boundary of $\Sigma$, and $\gamma_{i}$ is oriented by the flow-direction. This integer is called the multiplicity of the section $\Sigma$ at $\gamma_{i}$ and is the same for every component (see e.g. [24]).
[…]
If $\Sigma$ has genus $1$, since $\{\phi_{t}:M\to M\}_{t\in\mathbb{R}}$ is Anosov and $\mathring{\Sigma}$ is transverse to the $\phi$-orbits, then the center-stable and unstable foliations of the flow print on $\mathring{\Sigma}$ a pair of transverse foliations $\mathcal{F}^{\mathrm{s}}_{\Sigma}$, $\mathcal{F}^{\mathrm{u}}_{\Sigma}$, that extend onto foliations on $\hat{\Sigma}$ without singularities (since $\chi(\hat{\Sigma})=0$), that are respectively contracted and expanded by the action of $\hat{P}_{\Sigma}$. Thus, the homeomorphism $\hat{P}:\hat{\Sigma}\to\hat{\Sigma}$ is conjugated to a linear Anosov diffeomorphism on the torus given by a matrix $A\in\mathrm{SL}(2,\mathbb{Z})$ (cf. Lemma 1 at [15]). The Birkhoff section actually exhibits a path of Fried surgeries connecting $(\phi,M)$ with the suspension flow $(\phi^{A},M_{A})$:
$(\phi,M)]\xrightarrow[]{(\gamma_{1},-m_{1}),\dots,(\gamma_{k},-m_{k})}[(\phi^{A},M_{A})].$
@@
Given a periodic orbit $\alpha$ of the suspension flow $(\phi^{A},M_{A})$ define $\operatorname{\mathrm{per}}_{\mathbb{Z}}(\alpha)$ to equal the pairing $\alpha]\cdot[\mathbb{T}^{2}]$ of the homology class $\alpha]\in H_{1}(M_{A})$ with a generator $\mathbb{T}^{2}]$ of $H^{1}(M_{A};\mathbb{Z})$.
@@
Let $A\in\mathrm{SL}(2,\mathbb{Z})$ be a matrix with $\operatorname{\mathrm{tr}}(A)\geq 3$ satisfying that $A$ is conjugated to $A^{-1}$ in $\mathrm{GL}(2,\mathbb{Z})$. Then, given any periodic orbit $\gamma$ of the suspension flow $(\phi^{A},M_{A})$ there exists $m_{0}=m_{0}(\gamma)>0$ such that, for every integer $m$ with $|m|\geq m_{0}$, there are infinitely many Birkhoff sections $\iota:(\Sigma,\partial\Sigma)\to(M_{A},\Gamma)$ verifying that:
@@
      $p(\gamma,\Sigma)=\operatorname{\mathrm{per}}_{\mathbb{Z}}(\alpha)$, $p(\alpha,\Sigma)=\operatorname{\mathrm{per}}_{\mathbb{Z}}(\gamma)$,
@@ 2.5. Construction of the paths
Let’s prove Theorem A-A’ and Proposition B. For this, fix a matrix $A\in\mathrm{SL}(2,\mathbb{Z})$ with $\operatorname{\mathrm{tr}}(A)\geq 3$, choose some periodic orbit $\gamma$ of the suspension flow $\{\phi_{t}^{A}:M_{A}\to M_{A}\}_{t\in\mathbb{R}}$, let $m>0$ be an integer and define $(\psi,N)$ to be the flow obtained from the suspension $(\phi^{A},M_{A})$ by doing a Fried surgery along $\gamma$ with slope equal to $m$. We denote by $\Pi:\textsl{Per}(\phi^{A},M_{A})\to\textsl{Per}(\psi,N)$ the natural bijection between periodic orbits induced by this surgery, and denote $\delta=\Pi(\gamma)$.
@@
  • By Proposition 18, for every $m>0$ then $(\psi,N)$ is a skewed $\mathbb{R}$-covered Anosov flow, positively twisted;
@@
Setting $m\geq m_{0}$ we get a skewed $\mathbb{R}$-covered Anosov flow $\{\psi_{t}:N\to N\}_{t\in\mathbb{R}}$ on a hyperbolic 3-manifold. By Theorem 19, there exists an infinite set $\mathcal{FH}(\delta)$ of $\psi$-periodic orbits, each of them orientation preserving isotopic to $\delta$.
Figure 3. Surgery $(\phi^{A},M_{A})\xrightarrow[]{(\gamma,m),(\alpha,-m)}(\phi^{B},M_{A})$, $B\simeq A$ or $B\simeq A^{-1}$.
Consider one orbit $\beta\in\mathcal{FH}(\delta)$ different from $\delta$. Since $\delta$ and $\beta$ are equivalent as oriented knots, in particular there exists a homeomorphism on $N$ taking $\delta\mapsto\beta$ and preserving their orientations, so the surgery on $(\psi,N)$ with parameters $(\beta,-m)$ will produce an Anosov flow on a manifold homeomorphic to $M_{A}$. Recall that $M_{A}$ is a sol-manifold, therefore from the classification stated in Theorem 11 and Proposition 13 this flow must be orbitally equivalent to the suspension of either $A$ or $A^{-1}$. That is:
$(\phi^{A},M_{A})]\xmapsto[]{(\gamma,m)}[(\psi,N)]\xmapsto[]{(\beta,-m)}[(\phi^{B},M_{A})],\ \text{where}\ B\simeq A\ \text{or}\ B\simeq A^{-1}.$
If we assume in addition that $A$ is conjugated to $A^{-1}$ in $GL(2,\mathbb{Z})$ then the surgery described above is a closed loop in $\mathbf{G}_{\circ}$ with base point at the class of $(\phi^{A},M_{A})$. Define
  • $\alpha$ to be the preimage of $\beta$ under the surgery $(\phi^{A},M_{A})\xmapsto[]{(\gamma,m)}(\psi,N)$,
  • $\eta$ to be the image of $\delta$ under the surgery $(\psi,N)\xmapsto[]{(\beta,-m)}(\phi^{A},M_{A})$.
Let $\gamma^{\prime}$ be the image of $\beta$ under the surgery $(\psi,N)\xmapsto[]{(\beta,-m)}(\phi^{A},M_{A})$. Along the construction there is a sequence of homeomorphisms $M_{A}\setminus\gamma\to N\setminus\delta\to N\setminus\beta\to M_{A}\setminus\gamma^{\prime}$ so by Proposition 14 we have that, up to a self-orbital equivalence of ($\phi^{A},M_{A})$, the orbit $\gamma^{\prime}$ is equal to $\gamma$. We have codified paths:
(1) $[(\phi^{A},M_{A})]\xmapsto[]{(\gamma,m),(\eta,-m)}[(\phi^{A},M_{A})]\ \text{taking}\ \Pi:\eta\mapsto\gamma\ \text{and}\ \Pi:\gamma\mapsto\alpha$
(2) $[(\phi^{A},M_{A})]\xmapsto[]{(\gamma,m),(\alpha,-m)}[(\phi^{A},M_{A})]\ \text{taking}\ \Pi:\alpha\mapsto\gamma\ \text{and}\ \Pi:\gamma\mapsto\eta.$
The global transverse section of $\{\phi_{t}^{A}:M_{A}\to M_{A}\}_{t\in\mathbb{R}}$ is transformed by the surgery (1) into the desired Birkhoff section $\iota:(\Sigma,\partial\Sigma)\to(M_{A},\Gamma)$ for $(\phi^{A},M_{A})$, where $\Gamma=\{\gamma,\alpha\}$. In fact:
@@
  • It is direct that $\Sigma$ has $p(\alpha,\Sigma)=\operatorname{\mathrm{per}}_{\mathbb{Z}}(\gamma)$ boundary components along $\alpha$;
  • To determine $p(\gamma,\Sigma)$, since $\iota(\partial\Sigma)]=0\in H_{1}(M_{A})\simeq\mathbb{Z}$ we have the relation
    $0=p(\gamma,\Sigma)\cdot\operatorname{\mathrm{per}}_{\mathbb{Z}}(\gamma)\cdot m(\gamma,\Sigma)+p(\alpha,\Sigma)\cdot\operatorname{\mathrm{per}}_{\mathbb{Z}}(\alpha)\cdot m(\alpha,\Sigma),$
    from where we conclude that $p(\gamma,\Sigma)=\operatorname{\mathrm{per}}_{\mathbb{Z}}(\alpha)$.
Finally, the set $\mathcal{Q}(m,\gamma)$ consists in all periodic orbits $\alpha$ of $(\phi^{A},M_{A})$ such that $\beta=\Pi(\alpha)$ belongs to $\mathcal{FH}(\Pi(\gamma))$. ∎
@@
Proposition B follows from an estimation given in [2], on the total number of periodic orbits $\beta$ in $\mathcal{FH}(\delta)$ with period $\operatorname{\mathrm{per}}(\beta)\leq t$, for every $t\geq 0$.
@@
If $\{\psi_{t}:N\to N\}_{t\in\mathbb{R}}$ is a $\mathbb{R}$-covered Anosov flow in a hyperbolic manifold then there exist constants $C_{0},t_{0}>0$ such that, for every $\psi$-periodic orbit $\delta$, it is verified that
(3) $|\{\beta\in\mathcal{FH}(\delta):\operatorname{\mathrm{per}}(\beta)\leq t\}|\leq C_{0}\sqrt{t}\log(t),\ \forall\ t>t_{0}.$
@@
To prove Proposition B it is enough to show that there exist $C_{1},t_{1}>0$ such that
(4) $|\{\alpha\in\mathcal{Q}(\gamma,m):\operatorname{\mathrm{per}}(\alpha)\leq t\}|\leq C_{1}\sqrt{t}\log(t),\ \forall\ t>t_{1}.$
Because, in this case, it follows that
$\lim_{t\to\infty}\frac{\left|\left\{\alpha\in\mathcal{Q}(\gamma,m):\operatorname{\mathrm{per}}(\alpha)<t\right\}\right|}{\left|\left\{\alpha\in\textrm{Per}(\phi^{A},M_{A}):\operatorname{\mathrm{per}}(\alpha)<t\right\}\right|}\leq\lim_{t\to\infty}\frac{C_{1}\sqrt{t}\log(t)}{\kappa\cdot te^{\kappa t}}=0$
Here we use the fact that, for some constant $\kappa>0$, the number of periodic orbits with period less than $t$ is equivalent to $\kappa\cdot te^{\kappa t}$ as $t\to+\infty$, as provided by the Bowen-Margulis theorem (see, e.g. [19]).
To derive (4) from (3) we have to compare the periods of a $\phi^{A}$-periodic orbit $\alpha$ and the corresponding $\psi$-periodic orbit $\beta=\Pi(\alpha)$. To do this, let $T$ be a global transverse section of $(\phi^{A},M_{A})$. The surgery
$(\phi^{A},M_{A})\xrightarrow[]{(\gamma,m)}(\psi,N)$